(i) Design a voltage divider to produce an open-circuit output voltage of 4V from a voltage source of 10V. The current flowing in the circuit must be smaller than 1mA. (ii) If you connect a load of 1kΩ into the output of this voltage divider, what is the voltage drop across the load?


1
Expert's answer
2020-07-27T09:36:13-0400

(i) The circuit is shown on the picture below.



Here "U_{in} = 10V" and "U{out} = 4V".

The current which flows through this circuit is:


"I = \dfrac{U_{in}}{R_1 +R_2}"

Thus, the total resistance should be:


"R_1 + R_2 > \dfrac{U_{in}}{I} = \dfrac{10V}{1\times 10^{-3}A} = 10\times10^3\Omega = 10k\Omega"

Let's choose "R_1 = 10k\Omega". Then, according the voltage divider formula:

"U_{out} = U_{in}\dfrac{R_2}{R_1 + R_2}"

obtain the "R_2":


"R_2 = \dfrac{R_1U_{out}}{U_{in} - U_{out}} = \dfrac{10\cdot4}{10-4} \approx 6.67 k\Omega"

(ii) The circuit is shown on the picture below.



Let's conside resistors "R_2" and "R_{load}" as the lower leg of the voltage divider. Their total resistance will be (two resistors in parallel):


"R = \dfrac{R_2R_{load}}{R_2 + R_{load}} \approx0.87k\Omega"

Again, according the voltage divider formula:

"U_{out} = U_{in}\dfrac{R}{R_1 + R} = 10\dfrac{0.87}{10 + 0.87} = 0.8V"

Answer. (i) R1 = 10k, R2 = 0.87k, (ii) 0.8V.